# A universal neutral-atom quantum computer with individual optical addressing and non-destructive readout

### v2

@@ 4 Non-destructive State Selective Readout
Figure 4: Non-destructive state-selective readout. NDSSR signal distribution from 51862 images collected during 43 iterations of SPAM characterization. The turquoise-colored distribution describes the signals where atoms are expected to be in a bright state ($|F=4\rangle$) and the purple-colored distribution describes the signals for a dark state ($|F=3\rangle$). The samples for each distribution come from the deterministically loaded diagnostic site pair used for CZ gate fidelity benchmarking. Photo-electron counts for each site were calculated from the raw EMCCD camera analog-digital units using nominal EM gain and pre-amplifier gain. Each quadrant of the plane defined by NDSSR and post-NDSSR photo-electron counts contains two text labels showing total counts from $|F=4\rangle$ preparation and measurement experiment (turquoise) and counts from $|F=3\rangle$ preparation and measurement experiment (purple). a, Sampled probability distribution of photo-electron counts for bright and dark states. The dashed black line denotes the discrimination threshold that yields discrimination fidelity of 99.6(2)%. The shaded areas under the distributions on both sides of the discrimination threshold denote signals associated with state detection error. Depumping and atom loss during the NDSSR process and microwave pulse infidelity during state transfer from the stretched state to the bright qubit state give rise to bright state detection error of 3.1(7)%, which is reduced to 2.6(7)when atom loss is post-selected out. The dark state detection error of 1.6(5)is believed to be a combination of (partial) state preparation error and $F=3\rightarrow F=4$ repumping during imaging from residual leakage of repump laser intensity. The middle inset shows a single NDSSR measurement of a four-qubit GHZ state realized in the 24-qubit array. b, NDSSR and post-NDSSR occupancy readout photo-electron counts. The bright state atom loss rate is 1.0(4)and the dark state atom loss rate is 0.8(3)%. The atom loss rates and state detection errors and their uncertainties provided are the means and standard deviation of their measured values for the 43 experimental iterations that comprise the dataset.
@@
We implement the non-destructive state detection method demonstrated in Ref. [41] with a few critical improvements (see Methods 6.8). During the state detection sequence, qubit state $|1\rangle$ is optically pumped to the stretched state |$\ket{F=4,m_{F}=4}$, where we can detect fluorescence from the closed |$\ket{F=4,m_{F}=4}\rightarrow\ket{F^{\prime}=5,m_{F}^{\prime}=5}$ transition, while |$\ket{0}$ remains dark. By strategically alternating between illumination with trap light and readout light, we mitigate the deleterious effects of excited state light shift and dipole force fluctuations, and we reach a state-averaged loss rate of $0.86(25)\%$ while achieving bright-dark discrimination fidelity of 99.6(2)%with 6 ms of camera exposure time. The imaging performance and the state-discrimination fidelity of the readout scheme is analyzed from a large dataset of images collected during state preparation and measurement (SPAM) characterization experiments, each of which prepares all qubits to be in $|1\rangle$ (bright) or $|0\rangle$ (dark) (Fig. 4). The raw state-discrimination fidelity without state preparation or atom loss correction is 97.7(5)%, limited by bright state detection error of 3.1(7)and dark state detection error of 1.6(5)%. Note that state preparation error, which includes failure to optically pump into $|F=4,m_{F}=4\rangle$ during Raman sideband cooling and failure to coherently transfer $|F=4,m_{F}=4\rangle$ to $|F=4,m_{F}=0\rangle$ using microwave pulses (see Sec. 6.2), is distributed across both types of state detection errors.
The shot rate enhancement from repetitive non-destructive measurements is demonstrated by recording 8.2(2) Hz shot rate for a 15-qubit Greenberger–Horne–Zeilinger (GHZ) state preparation circuit executed within a 24-qubit array (see Sec. 6.8.3), compared to an estimated single-measurement shot rate of ${\sim}3$ Hz. These experiments were performed without Raman sideband cooling, the inclusion of which adds additional time to each measurement that cannot be amortized (see Methods Fig. 10). To the best of our knowledge, the only other demonstration of repeated non-destructive measurements of neutral-atom-based quantum circuits utilizing a universal gate set is found in Ref. [9]. This demonstration used an alkaline-earth species (strontium-88) with 10 ms of exposure time for non-destructive measurement of ancilla atoms with 3.5% atom loss rate.
@@ 5 Outlook
This work advances the state-of-the-art for individually addressed entangling gates in a neutral-atom platform, and we foresee a clear path towards the error levels required for fault tolerance. Control errors may be reduced by reaching lower atom temperatures, better stabilizing laser power and frequency, and using robust gate protocols[44, 45]. Straightforward changes such as increasing the intermediate-state detuning will immediately reduce scattering error. Moving forward, scattering error can be eliminated by transitioning to single-photon Rydberg excitation [38, 37], which also reduces the complexity of light shifts on atomic resonances. The path for scaling our qubit array is also well-established, with demonstrations[46] suggesting the feasibility of trapping $>10,000$ qubits in a single array, with the option to improve trapping lifetimes by operating at cryogenic temperatures[47]. The non-destructive readout we present here already enables state-of-the-art shot rate, and paths towards even faster readout are being developed[48, 49]. Finally, we note that our ability to achieve high ground-Rydberg coherence ($T_{2}^{*}=15~{}\mu\text{s}$) at $n=69$ is a unique advantage in the field, enabled by our electric field control. Since the Rydberg interaction scales with $n^{11}$, operating at higher $n$ unlocks greater connectivity, which is valuable for QEC protocols and applications[28].
For quantum computers to provide meaningful scientific and commercial value, they will likely need to support fault-tolerant operation on at least $\sim$100 logical qubits[50, 51, 5], and neutral-atom architectures are evolving rapidly towards that goal. This work demonstrates important new capabilities on that path, namely the ability to implement individually addressed high fidelity entangling gates and non-destructive state-selective readout, which collectively enable faster qubit operations and a broader set of QEC protocols. Ultimately, a compelling architecture for fast, fault-tolerant quantum computing is likely to integrate these techniques alongside other demonstrated neutral-atom capabilities, including dual-species platforms[52, 53], mid-circuit measurement[54, 55, 52, 53, 43, 7], and alternate addressing architectures[20, 56, 57, 7]. The advanced optical control techniques we demonstrate here also extend directly to alternative atomic species[20, 37, 58, 7], as well as multi-qubit ($N>2$) gates[59], which can further optimize logical operations[60].

### v3

@@ 4 Non-destructive State Selective Readout
Figure 4: Non-destructive state-selective readout. NDSSR signal distribution from 43 repetitions of SPAM characterization experiment. Each experiment executes two circuits - empty circuit and circuit with single $R(\pi,0)$ gate - with approximately 600 shots each, resulting in approximately 25800 camera frames for each circuit. The turquoise-colored distribution describes the signals where atoms are expected to be in a bright state ($|F=4\rangle$) and the purple-colored distribution describes the signals for a dark state ($|F=3\rangle$). The samples for each distribution come from the deterministically loaded diagnostic site pair used for CZ gate fidelity benchmarking. Photo-electron counts for each site were calculated from the raw EMCCD camera analog-digital units using nominal EM gain and pre-amplifier gain. Each quadrant of the plane defined by NDSSR and post-NDSSR photo-electron counts contains two text labels showing total counts from $|F=4\rangle$ preparation and measurement experiment (turquoise) and counts from $|F=3\rangle$ preparation and measurement experiment (purple). a, Sampled probability distribution of photo-electron counts for bright and dark states. The dashed black line denotes the discrimination threshold that yields discrimination fidelity of 99.6(2)%. The shaded areas under the distributions on both sides of the discrimination threshold denote signals associated with state detection error. Depumping and atom loss during the NDSSR process and microwave pulse infidelity during state transfer from the stretched state to the bright qubit state give rise to bright state detection error of 3.1(7)%, which is reduced to 2.6(7)when atom loss is post-selected out. The dark state detection error of 1.6(5)is believed to be a combination of (partial) state preparation error and $F=3\rightarrow F=4$ repumping during imaging from residual leakage of repump laser intensity. The middle inset shows a single NDSSR measurement of a four-qubit GHZ state realized in the 24-qubit array. b, NDSSR and post-NDSSR occupancy readout photo-electron counts. The bright state atom loss rate is 1.0(4)and the dark state atom loss rate is 0.8(3)%. The atom loss rates and state detection errors and their uncertainties provided are the means and standard deviation of their measured values for the 43 SPAM characterization experiments that comprise the dataset.
@@
We implement the non-destructive state detection method demonstrated in Ref. [47] with a few critical improvements (see Methods 6.8). During the state detection sequence, qubit state $|1\rangle$ is optically pumped to the stretched state |$\ket{F=4,m_{F}=4}$, where we can detect fluorescence from the closed |$\ket{F=4,m_{F}=4}\rightarrow\ket{F^{\prime}=5,m_{F}^{\prime}=5}$ transition, while |$\ket{0}$ remains dark. By strategically alternating between illumination with trap light and readout light, we mitigate the deleterious effects of excited state light shift and dipole force fluctuations, and we reach a state-averaged loss rate of $0.9(3)\%$ (1.0(4)for the bright state and 0.8(3)for the dark state) while achieving bright-dark discrimination fidelity of 99.6(2)with 6 ms of camera exposure time. The imaging performance and the state-discrimination fidelity of the readout scheme is analyzed from a large dataset of images of the diagnostic site pair used for CZ benchmarking analysis collected during state preparation and measurement (SPAM) characterization experiments, each of which prepares all qubits to be in $|1\rangle$ (bright) or $|0\rangle$ (dark) (Fig. 4). The raw state-discrimination fidelity without state preparation or atom loss correction is 97.7(5)%, limited by bright state detection error of 3.1(7)and dark state detection error of 1.6(5)%. Note that state preparation error, which includes failure to optically pump into $|F=4,m_{F}=4\rangle$ during Raman sideband cooling and failure to coherently transfer $|F=4,m_{F}=4\rangle$ to $|F=4,m_{F}=0\rangle$ using microwave pulses (see Sec. 6.2), is distributed across both types of state detection errors. The shot rate enhancement from repetitive non-destructive measurements depends on the quantum circuit being run, especially on the number of CZ gates as they can induce small atom loss. For an example circuit, we chose the preparation of a 15-qubit Greenberger–Horne–Zeilinger (GHZ) within a 24-qubit array, which contains 14 CZ gates. We recorded an average of 8.2(2) Hz shot rate for the example circuit compared to an estimated single-measurement shot rate of ${\sim}3$ Hz (see Sec. 6.8.3). These experiments were performed before the integration of Raman sideband cooling, which adds additional time to each measurement that cannot be amortized (see Methods Fig. 10).
@@ 5 Outlook
This work advances the state-of-the-art for individually addressed entangling gates in a neutral-atom platform, and we foresee a clear path towards the error levels required for fault tolerance. Control errors may be reduced by reaching lower atom temperatures, better stabilizing laser power and frequency, and using robust gate protocols[50, 51]. Straightforward changes such as increasing the intermediate-state detuning will immediately reduce scattering error. Moving forward, scattering error can be eliminated by transitioning to single-photon Rydberg excitation [43, 42], which also reduces the complexity of light shifts on atomic resonances. For larger systems, parallel gate operations using multiple beams is possible through multi-tone AOD drives and/or SLMs. The path for scaling our qubit array is also well-established, with demonstrations[52] suggesting the feasibility of trapping $>10^{4}$ qubits in a single array, with the option to improve trapping lifetimes by operating at cryogenic temperatures[53]. The non-destructive readout we present here already enables state-of-the-art shot rate, and paths towards even faster readout are being developed[54, 55]. Finally, we note that our ability to achieve high ground-Rydberg coherence ($T_{2}^{*}=15~{}\mu\text{s}$) at $n=69$ is a unique advantage in the field, enabled by our electric field control. Since the Rydberg interaction scales with $n^{11}$, operating at higher $n$ unlocks greater connectivity, which is valuable for QEC protocols and applications[32].
For quantum computers to provide meaningful scientific and commercial value, they will likely need to support fault-tolerant operation on at least $\sim$100 logical qubits[56, 57, 5], and neutral-atom architectures are evolving rapidly towards that goal. This work demonstrates important new capabilities on that path, namely the ability to implement individually addressed high fidelity entangling gates and non-destructive state-selective readout, which collectively enable faster qubit operations and a broader set of QEC protocols. Ultimately, a compelling architecture for fast, fault-tolerant quantum computing is likely to integrate these techniques alongside other demonstrated neutral-atom capabilities, including dual-species platforms[58, 59], mid-circuit measurement[60, 61, 58, 59, 49, 7], and alternate addressing architectures[24, 62, 63, 7]. In particular, future work should investigate the optimal combination of individual addressing, global beams, and mid-circuit rearrangement techniques to minimize circuit runtimes. For example, individual addressing can enable fast, parallel, short-range CZ gates, while motion unlocks long-range connectivity. And while a shared global beam can support parallel, identical CZ gates, individual addressing is preferable in the case of $R_{Z}(\phi)$ , to support parallel gates with distinct rotation angles. Additionally, individual addressing with tighter beams provides greater flexibility in achieving the desired laser intensity at significantly lower laser power, which is an important consideration for scaling and achieving higher-fidelity gates. The advanced optical control techniques we demonstrate here also extend directly to alternative atomic species[24, 42, 64, 7, 12, 13], as well as multi-qubit ($N>2$) gates[65], which can further optimize logical operations[66].